Minor Service Outage

About This Site

Welcome to The Signal Co's Network Status Page. This page provides updates of our network status. You can subscribe to Email and/or SMS updates to the network status and to individual components of our network.

We want you to know the cause of any issues should they occur. We see transparency as key to our success with our customers.

Backhaul, Interconnections and Systems Operational
Internet Capacity ? Operational
DNS Resolvers ? Operational
NTP Servers ? Operational
Buildings (ACT) Operational
90 days ago
99.96 % uptime
Today
470 Northbourne Ave, Dickson ? Operational
90 days ago
100.0 % uptime
Today
220 Northbourne Ave, Braddon ? Operational
90 days ago
100.0 % uptime
Today
13 Keltie St, Phillip ? Operational
90 days ago
99.98 % uptime
Today
120 Eastern Valley Way, Belconnen ? Operational
90 days ago
99.83 % uptime
Today
74 Chandler St, Belconnen Operational
90 days ago
100.0 % uptime
Today
20 Telita St, Bruce Operational
90 days ago
100.0 % uptime
Today
Towers (ACT) Operational
90 days ago
99.98 % uptime
Today
Isaacs Ridge (North) ? Operational
90 days ago
99.98 % uptime
Today
Isaacs Ridge (South) ? Operational
90 days ago
99.98 % uptime
Today
Tuggeranong Hill ? Operational
90 days ago
99.99 % uptime
Today
Strathnairn Operational
90 days ago
100.0 % uptime
Today
Towers (NSW) Operational
90 days ago
99.98 % uptime
Today
Sandra's View ? Operational
90 days ago
99.98 % uptime
Today
Klaus’s View ? Operational
90 days ago
99.98 % uptime
Today
Customer Interaction Platforms Operational
Phone System(s) Operational
Support Ticketing Operational
Customer Portal Operational
Website Operational
Credit Card Processing ? Operational
Status Page Accessibility ? Operational
Status Page SMS Notifications ? Operational
Status Page Email Notifications ? Operational
Third Party Systems used by TSC. Partial Outage
Cloudflare SYD - Sydney, NSW, Australia Operational
Cloudflare AKL - Auckland, New Zealand Operational
Cloudflare BNE - Brisbane, QLD, Australia Operational
Cloudflare MEL - Melbourne, VIC, Australia Partial Outage
Cloudflare PER - Perth, WA, Australia Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Major outage
Partial outage
No downtime recorded on this day.
had a major outage
had a partial outage
Past Incidents
Jul 22, 2019

No incidents reported today.

Jul 21, 2019

No incidents reported.

Jul 20, 2019

No incidents reported.

Jul 19, 2019

No incidents reported.

Jul 18, 2019

No incidents reported.

Jul 17, 2019

No incidents reported.

Jul 16, 2019

No incidents reported.

Jul 15, 2019

No incidents reported.

Jul 14, 2019

No incidents reported.

Jul 13, 2019

No incidents reported.

Jul 12, 2019

No incidents reported.

Jul 11, 2019
Completed - The scheduled maintenance has been completed.
Jul 11, 22:30 AEST
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Jul 9, 22:30 AEST
Scheduled - During the window we will be making network changes to migrate customer services throughout the network, This is hazardous work and may result in an unexpected outage due to human or software error.

Should this happen we apologise for any inconvenience caused.
Jul 9, 17:46 AEST
Completed - The scheduled maintenance has been completed.
Jul 11, 01:45 AEST
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Jul 10, 23:45 AEST
Scheduled - We will be undergoing scheduled maintenance during this time.
Jul 9, 17:42 AEST
Jul 10, 2019
Resolved - This incident has been resolved.
Jul 10, 23:59 AEST
Monitoring - We have implemented a majority of the upgrades and improvements, if your service is still speed limited, please power cycle (power off & power back on after 30 seconds) your rooftop unit and your router.

Your rooftop unit power supply can be found between your router and the wall socket.
Jul 10, 09:32 AEST
Update - Testing has completed, we are now working on migrating customer traffic.

Customers may experience a loss of data flow for approximately 10-30 mins during the works.

Works are being scheduled for 1am to 3am daily until complete in order to minimise disruption to services.
Jul 8, 19:49 AEST
Update - Optus have advised that the fibre provisioning is now complete.

Some minor provisioning tasks need to be completed before we can commence testing and go-live.

ETR Jul 10th.
Jul 5, 08:32 AEST
Update - We are waiting on an update from the upstream carrier in order to provide an update to the ETR.

At this point in time we suspect that it should all be good to go next week, however Optus haven't yet provided an updated RFS date beyond the RFS date of last Friday. We are pushing for this update as soon as possible.

For those technically interested, this is a WDM Layer 1 segment being delivered between the new data centre (and completely rebuilt core network) at NextDC in Bruce and our transmission hub at Lovett Tower in Woden which interconnects all of our wireless transmission sites.
Jun 28, 15:35 AEST
Update - Due to an embargo on changes by Optus their contractors have not yet been able to complete the fibre build for the new backhaul equipment to enable us to go live.

We understand and appreciate how frustrating this can be for our customers and we are working within the constraints of our third party providers and contractors to get this project completed ASAP.

Optus have been engaged to provide a detailed response to us on why this has been delayed, again.
Jun 28, 15:06 AEST
Update - ETR has been amended to 28th June
Jun 19, 21:40 AEST
Update - Stage 1. new equipment has arrived, our team will now configure and thoroughly test the new equipment configurations and performance.

Stage 2 will involve a traffic migration on a tower by tower basis.

We apologise for the delays, however, we believe that it is important to get this right the first time, we don't want to have any repeat issues.
Jun 19, 19:32 AEST
Identified - Our vendor has identified the issue and a fix is currently being tested before being implemented, ETR 24 Jun 19
Jun 13, 11:57 AEST
Update - We are continuing to investigate this issue. (Components Updated)
Jun 8, 12:40 AEST
Update - We are continuing to investigate this issue.
Jun 8, 12:39 AEST
Update - We are continuing to investigate this issue.
May 28, 15:06 AEST
Update - We are continuing to work with an equipment vendor to isolate this issue, There are upgrades in the pipeline that we are pushing elsewhere in the network that will alleviate/resolve this issue.

We have limited some customer speeds to DL ~10 Mbps UL ~10Mbps. These restrictions are temporary and will be removed as soon as possible.
May 21, 13:00 AEST
Investigating - We are investigating an issues that is transient and complex to pin-point, we apologise for any inconvenience caused. We will provide updates as they come to hand.
May 15, 19:48 AEST
Completed - The scheduled maintenance has been completed.
Jul 10, 01:25 AEST
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Jul 9, 23:45 AEST
Scheduled - We will be undergoing scheduled maintenance during this time.
Jul 9, 17:41 AEST
Completed - The scheduled maintenance has been completed.
Jul 10, 01:25 AEST
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Jul 9, 23:45 AEST
Scheduled - We will be undergoing scheduled maintenance during this time.
Jul 9, 17:40 AEST
Jul 9, 2019
Resolved - This incident has been resolved.
Jul 9, 12:00 AEST
Update - Root cause has been identified and resolved. Traffic will remain on backup paths until later this evening during a scheduled maintenance window.
Jul 9, 11:42 AEST
Monitoring - A fix has been implemented and we are monitoring the results.
Jul 9, 10:46 AEST
Investigating - We are currently investigating this issue.
Jul 9, 10:21 AEST